# Spin-glass dynamics: experiment, theory and simulation

### v1

@@ II.4 The elusive, but crucial, temperature chaos
Among the many interesting properties of spin glasses, the contextual measurement of both rejuvenation and memory effects has been an astonishing finding [162], which in turn asked for a theoretical understanding, eventually aided by numerical simulations of the EA model. These effects occur in the out-of-equilibrium regime and thus will be mainly covered in Sec. VI.2. Rejuvenation may, however, be related to an equilibrium effect, temperature chaos, which we briefly describe in this section. Furthermore, as we explain in Sec. II.7, temperature chaos is believed to be the main difficulty that impedes reaching thermal equilibrium even for a very small spin-glass sample.
The term chaos refers to the strong sensitivity of the spin-glass state —the one dominating the Gibbs-Boltzmann (GB) measure for $T<T_{\mathrm{g}}$— to the change of some parameter in the GB measure: either the temperature $T$ (temperature chaos), the couplings $\{J_{ij}\}$ (disorder chaos) or the external magnetic field $H$ (field chaos). These effects require the existence of a spin-glass phase, so they cannot take place in any dimensionality. When they are present, they have very different intensities, temperature chaos being by far the weakest. Indeed, for quite a long time the very existence of temperature chaos was controversial [see, e.g., 226, 225, 61, 217, 62, 184, 253, 259, 167, 246, 111], while field chaos and disorder chaos were evident.
In order to quantify the intensity of the chaotic effects, one needs to measure the statistical similarity between two typical configurations generated according to two GB measures differing in a small $\Delta T$ (for temperature chaos) or small $\Delta J$ (for disorder chaos). For example, for temperature chaos, one could measure the statistics of the following chaotic parameter
@@
If the sample under consideration is not chaotic, we expect the statistics (e.g., the variance) of overlap $q_{T_{1},T_{2}}$ between replicas at temperatures relatively close to one another to be very similar to those of $q_{T_{1},T_{1}}$ and $q_{T_{2},T_{2}}$, thus giving $X^{J}_{T_{1},T_{2}}\simeq 1$. On the contrary, under chaotic conditions we expect $q_{T_{1},T_{2}}$ to take the smallest possible value (which is zero if the external field is absent). In other words, in a chaotic sample, the configurations at temperatures $T_{1}$ and $T_{2}$ are typically very different (often maximally different).
@@
This is a very relevant physical phenomenon as it would imply that such a chaotic spin-glass sample would develop very different long-range order at temperatures $T_{1}$ and $T_{2}$. If chaos is present, it would be the most natural explanation for rejuvenation: the long-range order developed in the spin glass at the higher temperature $T_{2}$ would be replaced by a new long-range-ordered state when the sample is cooled down to the lower temperature $T_{1}$. How the first ordered state could be eventually recovered when the sample is heated back to $T_{2}$ is discussed in Sec. VI.2 and is the basis of the memory effect.
@@
The approaches to studying and quantifying chaotic effects in spin glasses have followed different paths: a scaling approach based on the droplet model (which is expected to be correct for low-enough dimensionalities); the solution to mean-field models, like the SK model (which is believed to be exact for high-enough dimensionalities); and the approach based on numerical simulations.
@@
According to the droplet theory, the low-temperature properties of a spin glass are controlled by a $T=0$ fixed point, and the large-scale excitations from the ground state (compact droplets with fractal dimension $D_{\text{s}}$) are controlled by an exponent $y$, which is positive (negative) if a spin-glass phase is present (absent). Using scaling arguments, Bray and Moore [71] showed that a perturbation of typical intensity $\Delta J$ to the couplings would destabilize the ground state on length scales larger than $\ell_{\text{c}}\approx(J/\Delta J)^{1/(D_{\text{s}}/2-y)}$, provided $D_{\text{s}}/2-y$ is positive.
@@
Moreover, they extended the argument to a temperature change $\Delta T$ (inside the spin-glass phase), finding again that a typical configuration would be modified on length scales larger than $\ell_{\text{c}}\sim(\Delta T)^{-1/1/(D_{\text{s}}/2-y)}$. So, for any temperature in the spin-glass phase, the overlap $q_{T_{1},T_{2}}$ would eventually trend to zero for large sizes. Within the droplet model, the picture one gets is very simple (probably too simple): the spatial correlation function between two replicas at different temperatures decays exponentially with a characteristic length scale $\ell_{\text{c}}$. Sample-to-sample fluctuations are not discussed, but we shall see that they play an important role in simulations.
@@
A complementary analytical approach can be developed for the mean-field SK model through the replica method [213]. In this case, it is more convenient to compute the free-energy cost $\Delta f(q_{\text{c}})=f(q_{\text{c}},T_{1},T_{2})-f(T_{1})-f(T_{2})$ to keep two replicas at different temperatures $T_{1}$ and $T_{2}$, with an overlap $q_{\text{c}}=q_{T_{1},T_{2}}\neq 0$. Indeed the probability distribution function of that overlap is
@@
This analytical result has convincingly clarified that spin glasses do have chaos in temperature (at least in the mean-field approximation), but the intensity of this phenomenon is very weak. This explains why it had not been observed for a long time in numerical simulation of the SK model [277, 263, 220, 64].
The elusiveness of temperature chaos is enhanced when we move away from the realm of mean-field models and study short-range systems. Numerical simulations of the EA model tell a complicated story, where sample-to-sample fluctuations play a key role and which builds on more recent equilibrium studies. Since these 3D results will be directly needed to understand some aspects of the non-equilibrium dynamics, we leave them for Sec. III.1.
@@ II.5 Advent of spin-glass single crystals
However, as Kasuya [166] has shown, electron scattering will change the long-range $\cos(2k_{\text{F}}\,r_{ij})/[k_{\text{F}}\,r_{ij}]^{3}$ dependence of the RKKY interaction to $r_{0}/[(r_{0}^{2}-r_{ij}^{2})^{2}+4r_{0}^{2}r_{ij}^{2}]$, where $k_{\text{F}}$ is the Fermi wave vector, $r_{ij}$ is the distance between spins $i$ and $j$, and $r_{0}$ is a characteristic scattering length that we take to be of the order of the crystallite size. The lack of an oscillatory character removes frustration, a requirement for spin-glass behavior (in addition to randomness), and thus effectively decouples the individual spin-glass grains, limiting the growth of the spin-glass correlation length. Adding to this difficulty is the distribution of grain sizes that must be taken into account for time-dependent properties.
All of the complications associated with poly-crystalline materials can be avoided by the use of single crystals. In that regard, a major step forward took place at the Materials Preparation Center (MPC) of the Ames Laboratory, U.S. Department of Energy, under the direction of D. L. Schlagel. She was able to fabricate two single-crystal CuMn boules with $T_{\mathrm{g}}=31.5$ K and 40.6 K, respectively. The details can be found in Appendix A of Zhai et al. [299]. A picture of the $T_{\mathrm{g}}=$ 31.5 K boule is reproduced in Fig. 15. Laue X-ray diffraction patterns were taken at various points along the length of the body of the same boule. A representative example is exhibited in Fig. 16, confirming that the majority of the body is one single grain. The spots make a clear pattern that is a qualitative measure of crystal quality. There is no evidence of twinning or strain in the crystal.
@@ II.6 Experimental extraction of $\xi(t,t_{\mathrm{w}};H)$
Experiments can measure the time dependence of the magnetization $M(t,t_{\mathrm{w}};H)$ at a fixed temperature. Nordblad et al. [228] introduced the derivative $S(t,t_{\mathrm{w}};H)$ [we rewrite Eq. (2) for the reader’s convenience]
@@
with the -- for TRM and $+$ for ZFC experiments. This quantity exhibits a maximum at a characteristic time $t_{\mathrm{w}}^{\text{eff}}$ that scales with the waiting time $t_{\mathrm{w}}$. Extraction of $\xi(t,t_{\mathrm{w}};H)$ from measurements of the time at which $S(t,t_{\mathrm{w}};H)$ has a maximum, yielding $t_{\mathrm{w}}^{\text{eff}}$, derives from Vincent et al. [282] using a free-energy “trap model” and Joh et al. [159] using a free-energy barrier model. These barriers or traps are created by the growth of $\xi(t,t_{\mathrm{w}};H)$ and they are invoked to account for the dynamical properties of spin glasses. Further, the deepest trap or the maximum barrier height controls the dynamics because of the exponential growth of occupancy with increasing barrier height. This can be seen from Fig. 14, and is a direct consequence of ultrametricity.
These works noted that the presence of a magnetic field reduces the depth of all the traps or, equivalently, the height of all the energy barriers, by the amount of energy associated with the presence of the magnetic field, or the “Zeeman energy” $E_{\text{Z}}$. They differed in their expressions for $E_{\text{Z}}$. For magnetic fields of laboratory strength the approach of Joh et al. [159] is the more appropriate.
Figure 17: $-M_{\text{TRM}}/H$ against measurement time $t$ (top), with the solid lines from the analytic fit, for CuMn 6 atfor various changes in magnetic field, and for $S(t,t_{\mathrm{w}};H)$ (bottom) calculated by taking the time derivative of the analytic fit (solid line) of $-M_{\text{TRM}}/H$ (top). Reproduced from Fig. 1 of Joh et al. [159].
@@
where $E_{\text{Z}},~{}N_{\text{s}}$, and $\chi_{\text{FC}}$ are defined below Eq. (10), the relationship between $N_{\text{s}}$ and the correlation length $\xi(t_{\mathrm{w}})$ exhibited in Eq. (11), and $\tau_{0}\approx\hbar/k_{\text{B}}T_{\mathrm{g}}$ is an “exchange time.” Thus, by measuring $t_{\mathrm{w}}^{\text{eff}}$ as a function of magnetic field $H$, one can extract $N_{\text{s}}$ and thence $\xi(t,t_{\mathrm{w}};H)$. This is the only vehicle one has available to determine $\xi(t,t_{\mathrm{w}};H)$ from experiment. The usual approaches of neutron diffraction or X-ray scattering measure only two-spin correlations. Unfortunately, for spin glasses, this averages to zero over sample dimensions. Instead, $\xi(t,t_{\mathrm{w}};H)$ derives from a four-spin correlation function [206, see Sec. IV.2].
The first extraction of $\xi(t,t_{\mathrm{w}};H)$ from experiments was performed by Joh et al. [159]. Using Eq. (16), they displayed the time dependence of $S(t,t_{\mathrm{w}};H)$ for different values of $H$ as displayed in Fig. 17. Following Eq. (17), they were able to plot the log $t_{\mathrm{w}}^{\text{eff}}$ vs $H^{2}$, as exhibited in Fig. 18. The break away from linearity with $H^{2}$ has now been ascribed to non-linear effects [231] rather than any breakdown of the relationships in Eq. (17).
Figure 18: A plot of log $t_{\mathrm{w}}^{\text{eff}}$ [equivalently $N_{\text{s}}\chi_{\text{FC}}H^{2}$ from Eq. (6)] vs $H^{2}$ for CuMn 6 at.($T/T_{\mathrm{g}}=0.83,t_{\mathrm{w}}=480$ s) (top) and for the insulating spin-glass CdCr1.7In0.3S4 ($T/T_{\mathrm{g}}=0.21,t_{\mathrm{w}}=3410$ s) at fixed $t_{\mathrm{w}}$ and $T$ (bottom). The dependence is linear for magnetic fields less than 250 and 45 G, respectively. Reproduced from Fig. 2 of Joh et al. [159].
The extraction of $\xi(t,t_{\mathrm{w}};T)$ follows immediately from Eq. (11) once $N_{\text{s}}\chi_{\text{FC}}H^{2}$ has been determined from plots such as Fig. 18. The combination of the use of this quantity, and the availability of single crystals discussed in Sec. II.5, has enabled the first direct comparisons between numerical and experimental measurements of the growth of spin-glass domains.
@@
II.7 Spin-glass simulations and dedicated computers

### v2

@@ II.4 The elusive, but crucial, temperature chaos
Among the many interesting properties of spin glasses, the contextual measurement of both rejuvenation and memory effects was an astonishing finding [191], which in turn asked for theoretical understanding, eventually aided by numerical simulations of the EA model. These effects occur in the out-of-equilibrium regime and will be mainly covered in Sec. VI.2. Rejuvenation may, however, be related to an equilibrium effect, temperature chaos, which we briefly describe in this section. Furthermore, as we explain in Sec. II.8, temperature chaos is believed to be the main difficulty that impedes reaching thermal equilibrium even for a very small spin-glass sample.
The term chaos refers to the strong sensitivity of the spin-glass state —the one dominating the Gibbs-Boltzmann (GB) measure for $T<T_{\mathrm{g}}$— to the change of some parameter in the GB measure: either the temperature $T$ (temperature chaos), the couplings $\{J_{ij}\}$ (disorder chaos) or the external magnetic field $H$ (field chaos). These effects require the existence of a spin-glass phase, so they cannot take place in any dimension less than the lower critical dimension $D_{\text{l}}\approx 2.5$ [145, 69, 236]. When they are present, they have very different intensities, temperature chaos being by far the weakest. Indeed, for quite a long time the very existence of temperature chaos was controversial [see, e.g., 281, 280, 64, 269, 65, 218, 317, 324, 196, 304, 128], while field chaos and disorder chaos were evident.
In order to quantify the intensity of the chaotic effects, one needs to measure the statistical similarity between two typical configurations generated according to two GB measures differing in a small $\Delta T$ (for temperature chaos) or small $\Delta J$ (for disorder chaos). For example, for temperature chaos, one could measure the statistics of the following chaoticity parameter
@@
If the sample under consideration is not chaotic, we expect the statistics (e.g., the variance) of overlap $q_{T_{1},T_{2}}$ between replicas at temperatures relatively close to one another to be very similar to those of $q_{T_{1},T_{1}}$ and $q_{T_{2},T_{2}}$, thus giving $X^{J}_{T_{1},T_{2}}\simeq 1$. Under chaotic conditions we expect $q_{T_{1},T_{2}}$ to take the smallest possible value (which is zero if the external field is absent). In other words, in a chaotic sample, the configurations at temperatures $T_{1}$ and $T_{2}$ are typically very different (often maximally different).
@@
This is a very relevant physical phenomenon. It would imply that such a chaotic spin-glass sample would develop very different long-range order at temperatures $T_{1}$ and $T_{2}$. If chaos is present, it would be the most natural explanation for rejuvenation: the long-range order developed in the spin glass at the higher temperature $T_{2}$ would be replaced by a new long-range-ordered state when the sample is cooled down to the lower temperature $T_{1}$. How the first ordered state could be recovered when the sample is heated back to $T_{2}$ is discussed in Sec. VI.2 and is the basis of the memory effect.
@@
The approaches to studying and quantifying chaotic effects in spin glasses have followed different paths: a scaling approach based on the droplet model (which is expected to be correct for low-enough dimensionalities); the solution to mean-field models, like the SK model (which is believed to be exact for high-enough spatial dimension); and the approach based on numerical simulations.
@@
According to the droplet theory, the low-temperature properties of a spin glass are controlled by a $T=0$ fixed point, and the large-scale excitations from the ground state (compact droplets with fractal dimension $D_{\text{s}}$) are controlled by an exponent $y$, which is positive (negative) if a spin-glass phase is present (absent). Using scaling arguments, Bray and Moore [76] showed that a perturbation of typical intensity $\Delta J$ to the couplings would destabilize the ground state on length scales larger than $\ell_{\text{c}}\sim(\Delta J)^{-1/[(D_{\text{s}}/2)-y]}$, provided $D_{\text{s}}/2-y$ is positive.
@@
Moreover, they extended the argument to a temperature change $\Delta T$ (inside the spin-glass phase), finding again that a typical configuration would be modified on length scales larger than $\ell_{\text{c}}\sim(\Delta T)^{-1/[(D_{\text{s}}/2)-y]}$. Hence, for any temperature in the spin-glass phase, the overlap $q_{T_{1},T_{2}}$ would eventually trend to zero for large sizes. Within the droplet model, the picture one gets is very simplistic (probably too simple): the spatial correlation function between two replicas at different temperatures decays exponentially with a characteristic length scale $\ell_{\text{c}}$. Sample-to-sample fluctuations are not discussed, but we shall see that they play an important role in simulations.
@@
A complementary analytical approach can be developed for the mean-field SK model through the replica method [257]. In this case, it is more convenient to compute the free-energy cost $\Delta f(q_{\text{c}})=f(q_{\text{c}},T_{1},T_{2})-f(T_{1})-f(T_{2})$ to keep two replicas at different temperatures $T_{1}$ and $T_{2}$, with an overlap $q_{\text{c}}=q_{T_{1},T_{2}}\neq 0$. The probability distribution function for that overlap is
@@
This analytical result has convincingly clarified that spin glasses do exhibit temperature chaos (at least in the mean-field approximation), but the intensity of this phenomenon is very weak. This explains why it had not been observed for a long time in numerical simulation of the SK model [347, 329, 275, 68].
The elusiveness of temperature chaos is enhanced when we move away from the realm of mean-field models and study short-range systems. Numerical simulations of the EA model tell a complicated story, where sample-to-sample fluctuations play a key role and which builds on more recent equilibrium studies. Because these 3D results will be directly needed to understand some aspects of nonequilibrium dynamics, we leave them for Sec. III.1.
@@ II.5 Advent of spin-glass single crystals
As Kasuya [195] has shown, electron scattering will change the long-range $\cos(2k_{\text{F}}\,r_{ij})/[k_{\text{F}}\,r_{ij}]^{3}$ dependence of the RKKY interaction to $r_{0}/[(r_{0}^{2}-r_{ij}^{2})^{2}+4r_{0}^{2}r_{ij}^{2}]$, where $k_{\text{F}}$ is the Fermi wave vector, $r_{ij}$ is the distance between spins $i$ and $j$, and $r_{0}$ is a characteristic scattering length that we take to be of the order of the crystallite size. The lack of an oscillatory character removes frustration, a requirement for spin-glass behavior (in addition to randomness), and thus effectively decouples the individual spin-glass grains, limiting the growth of the spin-glass correlation length. Adding to this difficulty is the distribution of grain sizes that must be taken into account for time-dependent properties.
All of the complications associated with poly-crystalline materials can be avoided by the use of single crystals. In that regard, a major step forward took place at the Materials Preparation Center (MPC) of the Ames Laboratory, U.S. Department of Energy, under the direction of D. L. Schlagel. She was able to fabricate two single-crystal CuMn boules with $T_{\mathrm{g}}=31.5$ K and 40.6 K, respectively. The details can be found in Appendix A of Zhai et al. [374]. A picture of the $T_{\mathrm{g}}=$ 31.5 K boule is reproduced in Fig. 16. Laue X-ray diffraction patterns were taken at various points along the length of the body of the same boule. A representative example is exhibited in Fig. 17, confirming that the majority of the body is a single grain. The spots make a clear pattern that is a qualitative measure of crystal quality. There is no evidence of twinning or strain in the crystal.
@@ II.6 Experimental extraction of $\xi(t,t_{\mathrm{w}};H)$
Experimentally, one needs a finite magnetization to make a measurement. This requires a magnetic field that can polarize the otherwise nonmagnetic quenched state. Bouchaud [70] introduced another piece of phenomenology, a Zeeman energy $E_{\text{Z}}$, which reduces each of the barrier heights —recall Sec. II.3— and is given by [188]
$E_{\text{Z}}=-M_{\text{FC}}H=-\chi_{\text{FC}}N_{\text{s}}H^{2},$ (13)
$N_{\text{s}}=\xi(t_{\mathrm{w}})^{D-\theta/2},$ (14)
where $\theta$ is the replicon exponent [35, 288, see also Eq. 25] and $D$ is the spatial dimension.
Specifically, one measures the time dependence of the magnetization $M(t,t_{\mathrm{w}};H)$ at a fixed temperature. Nordblad et al. [283] introduced the derivative $S(t,t_{\mathrm{w}};H)$ [we rewrite Eq. (2) for the reader’s convenience]
@@
with the - for TRM and ++ for ZFC experiments. This quantity exhibits a maximum at a characteristic time $t_{\mathrm{w}}^{\text{eff}}$ that scales with the waiting time $t_{\mathrm{w}}$. Extraction of $\xi(t,t_{\mathrm{w}};H)$ from measurements of the time at which $S(t,t_{\mathrm{w}};H)$ has a maximum, yielding $t_{\mathrm{w}}^{\text{eff}}$, derives from Vincent et al. [353] using a free-energy “trap model” and Joh et al. [188] using a free-energy barrier model. These barriers or traps are created by the growth of $\xi(t,t_{\mathrm{w}};H)$ and they are invoked to account for the dynamical properties of spin glasses. Further, the deepest trap or the maximum barrier height controls the dynamics because of the exponential growth of occupancy with increasing barrier height. This can be seen from Fig. 15, and is a direct consequence of ultrametricity.
These works note that the presence of a magnetic field reduces the depth of all the traps or, equivalently, the height of all the energy barriers, by the amount of energy associated with the presence of the magnetic field, or the “Zeeman energy” $E_{\text{Z}}$. They, however, differ in their expressions for $E_{\text{Z}}$. For magnetic fields of laboratory strength the approach of Joh et al. [188] is the more appropriate of the two.
@@
where $E_{\text{Z}},~N_{\text{s}}$, and $\chi_{\text{FC}}$ are defined below Eq. (13), the relationship between $N_{\text{s}}$ and the correlation length $\xi(t_{\mathrm{w}})$ exhibited in Eq. (14), and $\tau_{0}\approx\hbar/k_{\text{B}}T_{\mathrm{g}}$ is an “exchange time.” Thus, by measuring $t_{\mathrm{w}}^{\text{eff}}$ as a function of magnetic field $H$, one can extract $N_{\text{s}}$ and thence $\xi(t,t_{\mathrm{w}};H)$. This is the main vehicle one has available to determine $\xi(t,t_{\mathrm{w}};H)$ from experiment (see Sec. II.7 for possible alternatives). The usual approaches of neutron diffraction or X-ray scattering measure only two-spin correlations. Unfortunately, for spin glasses, this averages to zero over sample dimensions. Instead, $\xi(t,t_{\mathrm{w}};H)$ derives from a four-spin correlation function [245, see Sec. IV.2].
Combining the reduction of the energy barrier, Eq. (13), to the peak of the relaxation function $S(t,t_{\mathrm{w}};H)$ of Eq. (15) one obtains:
Because the replicon exponent $\theta$ depends on $\xi$ (see Fig. 13), the resolution of Eq. (17) is self-consistent in $\xi$. Zhai et al. [374] proposed an interpolation function for $\theta$ that holds for both the RSB and the droplet approach. [^2] Due to space constraints, we do not reproduce the full calculation here; it can be found in Appendix B of Zhai et al. [374]. It is commonplace to age the spin glass over a time $t_{\mathrm{w}}$, then probe its time dependent magnetization at a time $t$ after a magnetic field is turned on at $t_{\mathrm{w}}$, the zero-field-cooled (ZFC) magnetization. Alternatively, the system can be prepared in the presence of a magnetic field, at a time $t$ after which the field is reduced to zero at time $t_{\mathrm{w}}$, the thermoremanent magnetization (TRM); see Sec. I.2. The spin-glass correlation length is then denoted by $\xi(t,t_{\mathrm{w}};H)$.
Figure 18: $M_{\text{TRM}}/H$ against measurement time $t$ (top), with the solid lines from the analytic fit, for CuMn 6 at% for various changes in magnetic field, and for $S(t,t_{\mathrm{w}};H)$ (bottom) calculated by taking the time derivative of the analytic fit (solid line) of $M_{\text{TRM}}/H$ (top). Reproduced from Fig. 1 of Joh et al. [188].
Joh et al. [188] used Eq. (15) to display the time dependence of $S(t,t_{\mathrm{w}};H)$ for different values of $H$ as displayed in Fig. 18. Following Eq. (16), they were able to plot the log $t_{\mathrm{w}}^{\text{eff}}$ vs $H^{2}$, as exhibited in Fig. 19. The break away from linearity with $H^{2}$ has now been ascribed to nonlinear effects [288] rather than any breakdown of the relationships in Eq. (16).
Figure 19: A plot of log $t_{\mathrm{w}}^{\text{eff}}$ [equivalently $N_{\text{s}}\chi_{\text{FC}}H^{2}$ from Eq. (6)] vs $H^{2}$ for CuMn 6 at.% ($T/T_{\mathrm{g}}=0.83,t_{\mathrm{w}}=480$ s) (top) and for the insulating spin-glass CdCr1.7In0.3S4 ($T/T_{\mathrm{g}}=0.21,t_{\mathrm{w}}=3410$ s) at fixed $t_{\mathrm{w}}$ and $T$ (bottom). The dependence is linear for magnetic fields less than 250 and 45 G, respectively. Reproduced from Fig. 2 of Joh et al. [188].
The extraction of $\xi(t,t_{\mathrm{w}};H)$ follows immediately from Eq. (14) once $N_{\text{s}}\chi_{\text{FC}}H^{2}$ has been determined from plots such as Fig. 19. The combination of the use of this quantity, and the availability of single crystals discussed in Sec. II.5, has enabled the first direct comparisons between numerical and experimental measurements of the growth of spin-glass domains.
@@
II.7 Possible alternatives for extraction of $\xi(t,t_{\mathrm{w}};H)$
There have been other experimental methods suggested for the existence or extraction of $\xi(t,t_{\mathrm{w}};H)$ beyond the use of the magnetic-field dependence of the response function $S(t)$.
II.7.1 Nonlinear susceptibility
The nonlinear susceptibility in spin glasses was discovered by Miyako et al. [258]. It has been recently pointed out [67] that using the nonlinear susceptibility might be a nice way to access $\xi(t,t_{\mathrm{w}};H)$. In our context, one would study the magnetization $M(t,t_{\mathrm{w}})$ at several values of the external magnetic field $H$ and expand $M$ as
$M=\chi_{1}H+\frac{\chi_{3}}{3!}H^{2}+\frac{\chi_{5}}{5!}H^{5}+\ldots$ (18)
For sure, the study of nonlinear susceptibilities such as $\chi_{3}$ and $\chi_{5}$ is a classic topic in the spin-glass literature [66], particularly in an equilibrium context (i.e., for temperatures above $T_{\mathrm{g}}$). Furthermore, the experimental finding of scaling behavior for $\chi_{3}$ near $T_{\mathrm{g}}$ [165] played a major role in convincing the community that a thermodynamic phase transition takes place at $T_{\mathrm{g}}$, implying that spin-glass phenomenology is not merely a dynamic effect. Perhaps unsurprisingly, a similar approach has been followed in the study of supercooled liquids, a context where the physical origin of the glassy dynamics is still debated. In particular, the analogues of $\chi_{3}$ and $\chi_{5}$ have been measured in the dielectric polarizability of a supercooled liquid [55, 7, 8].
Given these precedents, it would be nice to extend this approach to the experimental study of spin glasses at $T<T_{\mathrm{g}}$ because it would bridge the study of magnetic disordered systems with that of supercooled liquids. A first step in this sense was taken in Baity-Jesi et al. [35], where $\chi_{3}(t,t_{\mathrm{w}})$ was investigated numerically. The following simple ansatz was found to approximately account for the simulation data at long times,
$\chi_{3}(t,t_{\mathrm{w}})=[\xi_{\mathrm{micro}}]^{D-\theta}G(t/t_{\mathrm{w}})\,$, (19)
where $G$ is a simple scaling function displaying a single maximum. The analogy between the above expression and the quantities employed in the analysis in Secs. V.2 and V.3 is striking.
II.7.2 Nonresonant hole burning
A series of papers has established nonresonant spectral hole burning (NSHB) as a vehicle for investigation of the nonexponential relaxation of glass-forming liquids and solids [92, 325, 90]. Chamberlin [91] applied a magnetic analogy of this technique to study the dynamical properties of spin glasses. His technique utilized a large-amplitude low-frequency pump oscillation, followed by a probe step, with a phase-cycling scheme. The technique requires that the sample absorb sufficient energy during the pump oscillation to modify its response spectrum, and that this modification persists long enough to be measured on laboratory time scales. Thus, NSHB can investigate linear response and thermal recovery.
Chamberlin [91] measured magnetic NSHB on a AuFe 5 at.% sample. He found that the bulk spectrum is heterogeneously broadened. Individual domains in the sample are heated, independent of one another. Their relaxation, on the time scale of the experiments, was also found to be independent of the other domains. As a function of recovery time, the monotonic return to equilibrium of the selected slow degrees of freedom shows that the domains are coupled to a common thermal bath, but with negligible direct coupling to the other slow degrees of freedom. This independence strongly suggests that the domains represent correlated regions with finite length scales. Thus, this study underpins the basis for the existence of the correlation length $\xi(t,t_{\mathrm{w}};H)$.
II.7.3 Mesoscopic electrical resistance noise measurements
Weissman and colleagues, in a series of papers [183, 361, 363, 362, 182, 9], proposed and showed experimentally that noise in the electrical resistance $R$ can provide a probe of spin dynamics in spin glasses because of universal conduction fluctuations (UCF) effects coupled to spin configurations. Their results for CuMn [182] displayed a minimum typical event size of $\approx 3\times 10^{4}$ spins at 13 K and $\approx 1.7\times~10^{5}$ spins at 23 K. The spectral density of $\delta R/R$ of “large samples” (with number of spins, $N_{\text{s}}>10^{12}$) rises sharply at a temperature $T_{\text{n}}$ near the spin-glass magnetic freezing temperature $T_{\mathrm{g}}$, and has a spectral shape sensitive to the magnetic field.
These results, when interpreted as the number of spins in a correlated volume, lead directly to the concept of a temperature dependent correlation length. Subsequent experimental results for CuMn [188] extracted the number of correlated spins as a function of the waiting time $t_{\mathrm{w}}$. The results for a temperature $T=28\text{ K}\approx 0.89T_{\mathrm{g}}$ varied from $1\times 10^{6}$ to $4\times 10^{6}$ (their Fig. 3), quite consistent with the $\delta R/R$ minimum estimates cited above for a typical event. The important point is that universal conduction fluctuations provided evidence for correlated spins in spin glasses, generating a minimum number that is consistent with later observations utilizing a completely different approach.
Further insight from UCF addressed alternate physical descriptions for spin-glass dynamics in Israeloff et al. [182]. To quote: “we find noise statistics in extreme disagreement with a dilute droplet or cluster model. We then determine a size-frequency scaling parameter which is inconsistent with a basic assumption of all droplet-like models [i.e., no sign is found of a size-frequency correlation, while the hierarchical picture fits in ‘surprising detail’]. This parameter and others are shown elsewhere [326, 131] to fit a hierarchical kinetic picture [228] inspired by the Parisi solution [256].” This finding is important because, for example, later extraction of the spin-glass coherence length $\xi(t,t_{\mathrm{w}};H)$ was unable to exhibit a difference between the prediction of the droplet and the hierarchical model [see Fig. 3 of 188].
II.8 Spin-glass simulations and dedicated computers